If a first order reaction undergoes 80 % completion in 30 minutes, then calculate the half life

Grade:12th pass

1 Answers

Arun
5 years ago
Order of reaction = 1Let the initial concentration be [A]0 = 100[A]t = 100 -80 = 20Rate law for first order rectionK = (2.303/t)log[A]0/[A]tSubstituting the values we getK = 0.05365Andt1/2 = 0.693/K = 12.91 minuts